9^(7-2)=1/243
please help,
at first i changed the left and right numbers into values that were similar
(3^2)^(7-2x) = 1/3^5
were would i go from here and explain the laws of indices needed to solve this problem

Answer by Alan3354(69443)

9^(7-2)=1/243
please help,
at first i changed the left and right numbers into values that were similar
(3^2)^(7-2x) = 1/3^5
-----------
The 1st "equation" has no x, then it appears.
(3^2)^(7-2x) = 1/3^5
(3^2)^(7-2x) = 3^(-5)
3^(14-4x) = 3^(-5)
14-4x = -5
x = 19/4